# Traversable wormholes inside anisotropic magnetized neutron stars: physical properties and potential observational imprints

### v1

@@ V Magnetized configuration of WH+NS systems
Approach 2 yields configurations with radii in the range of 0-$0.7$ km, as shown in Fig. 2. The corresponding MR curves exhibit significantly steeper trends compared to those obtained with Approach 1. For $B_{s}=10^{16}\,\mathrm{G}$, systems with low anisotropy parameter $h$ display the usual steep MR behavior, with only a mild flattening forming a shallow hump. In contrast, for anisotropic cases, a pronounced hump gradually develops as $h$ increases, while the steep branch becomes progressively shorter. In particular, for $h=2$, the maximum mass of the system no longer occurs along the steep branch of the MR curve, but instead is reached at the peak of the hump structure, although this peak remains relatively moderate and does not form an extreme maximum.
Figs. 8(a) and (c) show the stellar surface radius and the wormhole throat radius obtained using Approach 1 for $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$, respectively, while Figs. 8(b) and (d) display the corresponding results obtained using Approach 2. From Figs. 8(a), (b), and (c), we find that the wormhole throat radius curves for different values of the anisotropy parameter $h$ nearly coincide and exhibit almost identical slopes. This indicates that anisotropy primarily affects the range of the curves rather than their overall trend, consistent with the behavior observed in the non-magnetized configurations.
In contrast, for the configuration with $B_{s}=1\times 10^{16}\,\mathrm{G}$ within the framework of Approach 2, a different trend emerges: the wormhole throat radius curves exhibit noticeably different slopes. This behavior can be understood as a consequence of variations in the stellar surface radius $R_{s}$ in this case, which in turn influence the value of the throat radius $r_{0}$, since $r_{0}$ depends on $R_{s}$.
@@
Another important point to emphasize is that, in the magnetized configurations, Approach 2 yields systems with extremely small masses, even though their compactness values are comparable to those obtained using Approach 1. Objects with such properties are expected to be exceedingly difficult to detect observationally. Nevertheless, they remain of theoretical interest, and for this reason we still include them in our calculations of the surface redshift $z$ and the echo time $\tau$.
@@
Fig. 11 presents the echo time for each magnetized configuration obtained using Approach 1 and Approach 2. We find that a lower magnetic-field strength leads to a broader mass range in which gravitational-wave echoes may occur, corresponding to the mass interval where the system enters the ultracompact regime in terms of compactness. Within the adopted framework, Approach 1 yields a significantly wider mass range over which gravitational echoes can arise compared to Approach 2. We clearly observe that the trends of the echo-time curves in the magnetized configurations differ significantly from those of the non-magnetized case. In the presence of magnetic fields, the curves tend to become flatter, indicating that the echo duration varies only weakly across the considered mass range. This flattening suggests that the magnetic field suppresses the sensitivity of the echo time to variations in the ADM mass. Another important feature to be noted is the overall tendency of the echo-time curves in Approach 2, particularly for the configuration with $B_{s}=1\times 10^{16}\,\mathrm{G}$. In this case, the curves exhibit a decreasing trend with increasing mass. This behavior is especially pronounced for $h=1.5$ and $h=2$ within Approach 2. For $h=2$, the curve is no longer merely flat but clearly shows a decrease as the mass increases.
We observe that the magnetized configurations yield extremely large echo times, as shown in Fig. 11. Even for Approach 1 with $B_{s}=1\times 10^{15}\,\mathrm{G}$, the echo time can reach values of order $10^{12}\,\mathrm{s}$.On the other side, Approach 2 consistently produces smaller echo times. An intriguing feature emerges when comparing different magnetic-field strengths: the echo time for $B_{s}=1\times 10^{16}\,\mathrm{G}$ is systematically smaller than that for $B_{s}=1\times 10^{15}\,\mathrm{G}$. This behavior is counterintuitive, since one might expect stronger magnetic fields to further increase the echo time. For reference, the non-magnetized configurations yield echo times of order $10^{-1}$$10^{-2}\,\mathrm{ms}$, while the configuration with $B_{s}=1\times 10^{15}\,\mathrm{G}$ reaches the extremely large scale of $10^{12}\,\mathrm{s}$. Nevertheless, the configuration with $B_{s}=1\times 10^{16}\,\mathrm{G}$ produces a smaller echo time.
In order to clarify the origin of such nontrivial behavior, we further analyze the dependence of the echo time on the magnetic field strength $\mathcal{B}$. To address this, we consider a toy model in which the magnetic field is taken to be constant from $r_{0}$ to $R_{s}$, and the pressure is assumed to be isotropic, such that the conservation law yields
@@
In Fig. 12, we observe that the echo time tends to remain nearly constant over a wide range of the magnetic field strength $\mathcal{B}$, taking values of the order of $10^{-1}$ ms. This magnitude is comparable to that obtained in the non-magnetized configuration. So it is clear that the echo time at $\mathcal{B}\rightarrow 0$ is consistent with the echo time shown by Fig. 5. The echo time exhibits a flat behavior up to $\mathcal{B}\sim 10^{16}$ G, beyond which it drops sharply. This qualitative trend is consistent with the behavior found in the magnetized configurations shown in Fig. 11.
The appearance of extremely large echo times, reaching very high orders of magnitude in magnetized configurations, can be attributed to several factors entering the echo-time integral. In particular, the complicated functional form of the metric function $e^{2\nu(r)}$ plays an important role. Moreover, these large values are also related to the fact that the mass and radius are not kept fixed in those configurations, and can become significantly larger than the corresponding values in Set 1 and Set 2.
Physically, this implies that the gravitational-wave signal propagates over a much longer effective distance under an extremely strong gravitational field in the magnetized configurations. Consequently, it is not surprising that the resulting echo time becomes very large.
Nevertheless, even within this simplified toy-model analysis, the echo-time profile as a function of the magnetic field provides useful insight. In particular, it clearly indicates that the echo time may decrease for strong surface magnetic fields around $B_{s}=10^{16}$ G, offering a plausible explanation for the behavior we see in the magnetized configurations presented by Fig. 11.
VI Conclusion
In this study, we have constructed the formalism of anisotropic WH+NS systems, both in the absence and in the presence of magnetic fields. In general, the underlying model admits ghosts; however, these ghosts can be eliminated by imposing appropriate constraints through the introduction of Lagrange multiplier fields in the extended action. We find that neither anisotropy nor magnetic fields affect the traversability of the wormhole, in the sense that both ingredients do not modify the violation of the NEC in the vicinity of the wormhole throat.
We investigate the physical properties of the resulting configurations, including the ADM mass, stellar radius, wormhole throat radius, and compactness. In addition, we analyze potential observational imprints, namely the surface redshift and the gravitational-wave echo time. Throughout our analysis, we adopt two complementary approaches, following the methodology introduced in Ref. [38]. In Approach 1, for a given central energy density $\rho_{c}$, the stellar radius is assumed to coincide with that of an ordinary NS obtained at the same $\rho_{c}$. In Approach 2, for the same $\rho_{c}$, the total mass of the WH+NS system is taken to be equal to that of an ordinary NS computed at the same central energy density.
@@ VI Conclusion
Furthermore, we find that both anisotropy and magnetic fields have only a minor influence on the size of the wormhole throat itself. Their primary effect is instead on the allowed range of the throat radius, which extends to larger values as either the anisotropy parameter $h$ or the surface magnetic field increases. This behavior can be traced back to the fact that larger anisotropy and stronger magnetic fields can lead to more massive configurations.
In the context of the surface redshift, we find that all configurations considered in this work can exceed values of $z\gtrsim 1.5$, which are significantly larger than those typically expected for ordinary neutron stars. Such large surface redshift values therefore provide a promising potential observational imprint of WH+NS systems. Furthermore, for a fixed configuration and at a given mass, larger values of the anisotropy parameter $h$ systematically lead to smaller surface redshift values.
The echo time for non-magnetized configurations is of the order of $10^{-2}$$10^{-1}$ ms. In contrast, for magnetized configurations the echo time can become extremely large. For a surface magnetic field of $B_{s}=1\times 10^{15}$ G, the maximum echo time reaches values of the order of $10^{12}$ s, whereas for $B_{s}=1\times 10^{16}$ G the maximum value is reduced to approximately $10^{10}$ s.
In addition, we construct a formulation of the echo time as an explicit function of the magnetic field, assuming a constant magnetic field throughout the system. The resulting behavior exhibits a nearly flat profile from $B\simeq 0$ up to $B\sim 10^{15}$ G, followed by a sharp decrease near $B\sim 10^{16}$ G. This behavior provides a clear explanation for why the echo time associated with a surface magnetic field of $B_{s}=1\times 10^{16}$ G becomes smaller than that obtained for $B_{s}=1\times 10^{15}$ G.
@@ Acknowledgments
MLP gratefully acknowledges Byon Jayawiguna for precious discussions. MLP is also financially supported by Indonesia Endowment Fund for Education Agency (LPDP). HLP was funded by the postdoctoral program in BRIN (National Research and Innovation Agency), Indonesia. MFARS is supported by the Second Century Fund (C2F) and C2F research abroad scholarship. We would like to thank the members of the Theoretical Physics Groups at Institut Teknologi Bandung for the discussions, insights, and hospitality.
@@
In this case, Eq. (62) gives
@@
Note that the condition shown by Eq. (72) is also satisfied in this case. Eq. (65) is now reduced and becomes

### v3

@@ 5 Magnetized configuration of WH+NS systems
Approach 2 yields configurations with radii between 0-$0.7$ km, as shown in Fig. 7(b) and (d). The corresponding MR curves exhibit significantly steeper trends compared to those obtained with Approach 1. For $B_{s}=10^{16}\,\mathrm{G}$, systems with low anisotropy parameter $h$ display the usual steep MR behavior, with only a mild flattening forming a shallow hump. In contrast, for anisotropic cases, a pronounced hump gradually develops as $h$ increases, while the steep branch becomes progressively shorter. In particular, for $h=2$, the maximum mass of the system no longer occurs along the steep branch of the MR curve, but instead is reached at the peak of the hump structure, although this peak remains relatively moderate and does not form an extreme maximum.
The behavior observed in Fig. 7 can be understood by recalling how the reference scales are fixed in the two approaches. In both cases, the calibration is performed with respect to the corresponding pure NS solutions. In Approach 1, the surface radius of the WH+NS system is fixed to coincide with the radius of the pure NS, while in Approach 2 the total ADM mass is fixed to the mass of the pure NS at the same central density.
In the non-magnetized configuration, the masses and radii of pure NSs lie within a moderate and physically familiar range. The obtain masses and radii of the WH+NS systems also lie within moderate range, and comparable to MR relation of the pure NS, as illustrated in Fig. 1.
The inclusion of a magnetic field qualitatively alters this picture. In magnetized configurations, the wormhole throat radius $r_{0}$, and therefore the ADM mass $M=r_{0}/2$, explicitly depend on the magnetic field parameters. At large radii, where the mass of a pure magnetized NS becomes very small, a WH+NS configuration constructed using Approach 1 necessarily develops an extremely large mass as the consequence of the Eq. (77). This leads to the appearance of very massive WH+NS systems, as shown in Figs. 7(a) and (c).
On the other hand, when using Approach 2, the ADM mass of the WH+NS system is constrained to remain equal to that of the pure magnetized NS. Under this condition, the relation $M=r_{0}/2$ forces the throat radius and the surface radius to shrink to very small values. The extremely small masses and radii observed in Figs. 7(b) and (d) therefore reflect the non-trivial interplay between the magnetic field and the imposed mass constraint, rather than any numerical instability.
Overall, the magnetic field gives rise to WH+NS configurations with MR properties that differ qualitatively from those of the non-magnetized case: Approach 1 naturally leads to extremely massive systems, whereas Approach 2 produces configurations with extremely small radii.
Fig. 8 shows the wormhole throat radius $r_{0}$ and surface radius $R_{s}$ as the functions of the ADM mass $M$ for different values of the anisotropy parameter $h$ and the different surface magnetic field $B_{s}$, within two construction approaches considered in this work. In panels (a) and (c), corresponding to Approach 1, we can obvioulsy see that the throat-radius curves coincide. In panel (b), corresponding to Approach 2, the solutions exist only within distinct mass intervals. As a result, the throat radius curves for $h=1.5$ and $h=2$ appear separated. This behavior does not indicate a modification of the functional relation $r_{0}(M)$, but rather reflects the restricted mass ranges over which the corresponding solutions are physically admissible. A different behavior is seen in panel (d), which also corresponds to Approach 2. Here, we can see that the curves of the throat radius are separated to several disconnected mass intervals. Within each interval the relation $r_{0}=2M$ is still satisfied, but the allowed ranges of the throat radius differ, which explains the clearly separated curves shown in the figure. It is important to note that anisotropy does not modify the geometric relation between $r_{0}$ and $M$, but it does affect the domain of existence of WH+NS solutions, with a particularly strong impact in Approach 2.
@@
Another important point to emphasize is that, in the magnetized configurations, as mentioned before, Approach 2 yields systems with extremely small masses, even though their compactness values are comparable to those obtained using Approach 1. Objects with such properties are expected to be exceedingly difficult to detect observationally. Thus, Approach 2 is not relevant for obtaining physically detectable objects (at least at the present time). Nevertheless, they remain of theoretical interest, and for this reason we still include them in our calculations of the surface redshift $z$ and the echo time $\tau$.
@@
Fig. 11 presents the echo time for each magnetized configuration obtained using Approach 1 and Approach 2. For Approach 1, both configurations with $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$ yield echo times of order $10^{-1}$ ms, which is the same order of magnitude as in the non-magnetized case. However, the echo times in the magnetized configurations are systematically larger than those obtained without magnetic fields. In the non-magnetized configuration, the longest echo time considered corresponds to $h=2$ and remains below $0.4$ ms. By contrast, for $B_{s}=1\times 10^{15}\,\mathrm{G}$ the maximum echo time at $h=2$ exceeds $0.6$ ms, while for $B_{s}=1\times 10^{16}\,\mathrm{G}$ it can exceed $0.9$ ms. This behavior can be attributed to the larger surface radii induced by the magnetic field, which increase the characteristic propagation distance of the echo signals.
On the other hand, Approach 2, shown in Figs. 11(b) and (d), yields extremely short echo times, of order $10^{-2}\,\mu\mathrm{s}$ for $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $10^{-1}\,\mu\mathrm{s}$ for $B_{s}=1\times 10^{16}\,\mathrm{G}$. This behavior can be understood as a consequence of the very small separation between the wormhole throat radius $r_{0}$ and the stellar surface $R_{s}$, which confines the echo propagation to a very short radial distance. We note that such short echo times would be extremely challenging to detect. Nevertheless, as emphasized earlier, Approach 2 is primarily of theoretical interest and is included to illustrate the range of possible behaviors allowed by the model.
In general, a direct comparison between the systems with $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$ is not straightforward, since the mass ranges over which echo signals exist differ between the two cases. An exception occurs for the $h=1$ configuration in Approach 1, where the admissible mass intervals overlap in the range $M\simeq 3.90$$4.17\,M_{\odot}$, allowing a meaningful comparison. The comparison is shown by Table 2.
Table 2: Echo time for magnetized WH+NS systems with $h=1$ for different masses. The comparison is shown for surface magnetic fields $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$.
$M\,(M_{\odot})$ $\tau$ (ms)
$B_{s}=1\times 10^{15}\,\mathrm{G}$ $B_{s}=1\times 10^{16}\,\mathrm{G}$
3.90 0.24 0.20
3.97 0.25 0.21
4.03 0.26 0.22
4.12 0.27 0.23
4.17 0.29 0.24
Since we obtain the echo times in magnetized configurations that are larger than those in the non-magnetized case, one might expect the echo time to increase monotonically with the magnetic-field strength. However, Table 2 shows that, within the mass range where a direct comparison is possible, the systems with $B_{s}=1\times 10^{16}\,\mathrm{G}$ exhibit shorter echo times than those with $B_{s}=1\times 10^{15}\,\mathrm{G}$. In order to clarify the origin of such nontrivial behavior, we further analyze the dependence of the echo time on the magnetic field strength $\mathcal{B}$. To address this, we consider a toy model in which the magnetic field is taken to be constant from $r_{0}$ to $R_{s}$, and the pressure is assumed to be isotropic, such that the conservation law yields
@@
In Fig. 12, we observe that the echo time tends to remain nearly constant over a wide range of the magnetic field strength $\mathcal{B}$, taking values of the order of $10^{-1}$ ms. This magnitude is comparable to that obtained in the non-magnetized configuration. So it is clear that the echo time at $\mathcal{B}\rightarrow 0$ is consistent with the echo time shown by Fig. 5. The echo time exhibits a flat behavior up to $\mathcal{B}\sim 10^{16}$ G, beyond which it drops sharply.
Within this simplified toy-model analysis, the echo-time profile as a function of the magnetic field provides useful insight. In particular, it clearly indicates that the echo time may decrease for strong surface magnetic fields around $B_{s}=10^{16}$ G, offering a plausible explanation for the behavior we see in the magnetized configurations presented by Table 2.
6 Conclusion
In this study, we have constructed the formalism of anisotropic WH+NS systems, both in the absence and in the presence of magnetic fields. In general, the underlying model admits ghosts; however, these ghosts can be eliminated by imposing appropriate constraints through the introduction of Lagrange multiplier fields in the extended action. We find that the wormhole remains traversable irrespective of whether fluid anisotropy and/or magnetic fields are present, in the sense that the NEC is violated in the vicinity of the wormhole throat both with and without these effects.
We investigate the physical properties of the resulting configurations, including the ADM mass, stellar radius, wormhole throat radius, and compactness. In addition, we analyze potential observational imprints, namely the surface redshift and the gravitational-wave echo time. Throughout our analysis, we adopt two complementary approaches, following the methodology introduced in Ref. Nojiri2024 . In Approach 1, for a given central energy density $\rho_{c}$, the stellar radius is assumed to coincide with that of an ordinary NS obtained at the same $\rho_{c}$. In Approach 2, for the same $\rho_{c}$, the total mass of the WH+NS system is taken to be equal to that of an ordinary NS computed at the same central energy density.
@@ 6 Conclusion
In the context of the surface redshift, we find that all configurations considered in this work can exceed values of $z1.5$, which are significantly larger than those typically expected for ordinary neutron stars. Such large surface redshift values therefore provide a promising potential observational imprint of WH+NS systems. Furthermore, for a fixed configuration and at a given mass, larger values of the anisotropy parameter $h$ systematically lead to smaller surface redshift values.
The echo time for non-magnetized configurations is of the order of $10^{-2}$$10^{-1}$ ms. For magnetized configurations, the echo time in Approach 1 is of order $10^{-1}\,\mathrm{ms}$ and is larger than that of the corresponding non-magnetized configurations. In contrast, Approach 2 yields much shorter echo times for magnetized systems, of order $10^{-2}$$10^{-1}\,\mu\mathrm{s}$.
In addition, we construct a formulation of the echo time as an explicit function of the magnetic field, assuming a constant magnetic field throughout the system. The resulting behavior exhibits a flat behavior from $B\simeq 0$ up to $B\sim 10^{15}$ G, followed by a sharp decrease near $B\sim 10^{16}$ G. This trend explains why, within the mass range where a direct comparison is possible, the echo times associated with a surface magnetic field $B_{s}=1\times 10^{16}\,\mathrm{G}$ are smaller than those obtained for $B_{s}=1\times 10^{15}\,\mathrm{G}$.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, M. Lawrence Pattersons used ChatGPT (OpenAI) in order to support brainstorming and organization of some theoretical ideas of this work, and to improve the language and readability of the manuscript. All physical interpretations, model constructions, and final manuscript preparation were performed by the authors, who take full responsibility for the content. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the paper.
@@ Acknowledgments
MLP gratefully acknowledges Byon Jayawiguna for precious discussions on the echo time. MLP is also financially supported by Indonesia Endowment Fund for Education Agency (LPDP). HLP was funded by the postdoctoral program in BRIN (National Research and Innovation Agency), Indonesia. MFARS is supported by the Second Century Fund (C2F) and C2F research abroad scholarship. We would like to thank the members of the Theoretical High Energy Physics Group at Institut Teknologi Bandung for the discussions, insights, and hospitality.
@@
In this case, Eq. (LABEL:numagnetized) gives
@@
Note that the condition shown by Eq. (82) is also satisfied in this case. Eq. (LABEL:derivativenu) is now reduced and becomes